# Implications of DESI for Dark Matter & Cosmic Birefringence

### v1

@@ 3.1 Dark matter in polynomial potential
[…]
@@
where we consider $n_{\rm eq}^{\rm DM}(T)=\frac{T}{2\pi^{2}}\,m_{\psi}^{2}\,K_{1}\left(m_{\psi}/T\right)$ to be the equilibrium DM number density, such that it provides a conservative limit on $T_{\text{rh}}$. For our choice of parameters, Eq. (41) is always satisfied.
@@ 3.2 Dark matter in axion-type potential
    Dark freeze-out: The dark baryons $\psi$ can undergo freeze-out within the DS via, for example, $\bar{\psi}\psi\to\pi_{\rm D}\,\pi_{\rm D}$, where $\pi_{\rm D}$ are the dark pions. The dark quarks $Q,\bar{Q},{\mathbf{u,d}}$ can remain in equilibrium with each other via the mediation of dark gluons. Therefore, a notion of dark temperature $T_{\rm D}$ can be attributed. Despite $Q,\bar{Q}$ carrying hypercharge quantum numbers, and in the presence of dimension-5 EM coupling and effective operators relevant for $Q,\bar{Q}$ decay as mentioned in subsection 2.3, it is easy to ensure that the DS never comes in thermal equilibrium with the visible sector in the entire thermal cosmological history due to very large $f$. As long as one satisfies $T_{\rm rh}<f$, the heavy dark quarks $Q,\bar{Q}$ never become massless in the cosmological evolution. For example, consider the scattering $\gamma Q\to\gamma Q$ in the early Universe. It is sufficient to satisfy that the decay rate for this process is less than the Hubble rate at the reheating temperature. This is satisfied when
@@
    obtained by comparing the interaction rate $n_{\rm eq}\,\langle\sigma\cdot v\rangle$ with the Hubble rate in a radiation dominated background, where we consider $\langle\sigma\cdot v\rangle\simeq C_{Y}^{4}\,\pi\,\alpha_{1}^{2}/m_{Q}^{2}$ and $n_{\rm eq}=\left(2\,\zeta(3)/\pi^{2}\right)\,T^{3}$ is the equilibrium number density of photons. The condition in Eq. (42) is easily satisfied with the chosen benchmark parameters in subsection 2.3. Similarly, we have checked that it is straightforward to satisfy the out-of-equilibrium condition for the dimension-5 EM coupling and the effective operators relevant for $Q,\bar{Q}$ decay. In particular, if one considers the operator $\langle\Phi\rangle^{*}/\Lambda_{d}^{3}\bar{\mathbf{u}}_{R}Qe_{R}e_{R}$, then the DS never comes in thermal equilibrium with the SM, provided
@@
    This points towards the dark QCD with dynamical scale $\Lambda_{\rm dQCD}\sim 10$ TeV, and therefore, the DM mass is $m_{\psi}\sim{\cal O}(50)$ TeV for $N_{D}=3$ as determined from Eq. (29). In this case, the DM behaves collisionless even at the dwarf scale as can be seen from Eq. (27).
@@
    Let us now turn to the question of stable DS relics other than the DM. For the current model, there are three dark pions $\pi_{\rm D}$, with mass $m_{\pi_{\rm D}}\simeq\sqrt{m_{\mathbf{q}}\Lambda_{\rm dQCD}}$, and can be determined from Eq. (26) for a given $\Lambda_{\rm dQCD}$ and $\varepsilon$. Clearly, for $\Lambda\simeq 2$ meV, as required by the DE energy density, and $\Lambda_{\rm dQCD}\simeq 10$ TeV, as required by the dark freeze-out mechanism, $m_{\pi_{\rm D}}\lesssim 10^{-14}$ eV for $\varepsilon\gtrsim 10^{-10}$. Further, in the absence of dark electromagnetism and weak interaction, the pions are stable. They also do not decay to the DE axion $\varphi$ due to very suppressed mixing $\simeq\Lambda_{\rm dQCD}/(4\pi f)\ll 1$. Hence, they behave as stable dark radiation and can be constrained by extra number of relativistic degrees of freedom $\Delta N_{\rm eff}$. If one assumes that the entropy is separately conserved in the visible and in the DS, then it is possible to write, $\xi\propto\left[g_{\star s}(T)/g_{\star s}^{\rm D}(T_{\rm D})\right]^{1/3}$, where $g_{\star s}$ and $g_{\star s}^{\rm D}$ are the number of relativistic degrees of freedom for entropy in the visible and the DS, respectively. As discussed in Cheung:2010gj , it is possible to constrain $\xi$ from the measurement of $\Delta N_{\text{eff}}=\frac{4}{7}\,g_{\star}^{\rm D}(T_{\rm BBN})\,\xi(T_{\rm BBN})^{4}$, that takes care of the number of extra relativistic degrees of freedom beyond the SM at the time of BBN. Considering $\Delta N_{\text{eff}}\lesssim 0.34$ Planck:2018vyg , one finds, $\xi(T_{\rm BBN})\lesssim 0.67$, for $g_{\star}^{D}(T_{\rm BBN})=3$, as relevant for the 3 dark pions. It is also possible to introduce portal operators for the dark pions to decay to SM particles, and then both the visible and DS can be in thermal equilibrium, and can introduce laboratory probe channels for the DM Fujikura:2024jto . We do not pursue this possibility here.
    In passing, we mention that in the present scenario, it is possible to have dark glueballs generated in the thermal bath. However, they will be unstable and decay into the dark pions, given that the dark pion mass is tiny. Further, if the confining phase transition of this dark QCD sector is first-order, which is expected for large $N_{\rm D}$, it could be associated with gravitational waves (GWs). It is expected that the scale $\Lambda_{\rm dQCD}$ would be responsible for setting the reheating temperature scale after this phase transition. Therefore, the gravitational waves would have a peak frequency in the range Santos:2022hlx ; Fujikura:2023lkn
@@ 4 Conclusions
In this work, we have explored a consistent and minimal framework for an interacting DE-DM scenario that successfully explains the phantom crossing behavior of the evolving DE equation-of-state parameter, as indicated by recent DESI observations, without inducing any pathological features. We have introduced a dimension-5 effective operator involving the DE scalar and the SM photons, which accounts for the observed non-zero ICB angle. To explain the evolving DE behavior, we consider two well-motivated potentials for the DE scalar field: a polynomial potential, representative of tracker behavior in conventional quintessence models, and an axion-like potential naturally arising in strongly-coupled QCD-like dark sectors. We have provided a UV complete model for the latter. Furthermore, we have demonstrated how the DM can minimally obtain its relic density. Interestingly, the polynomial potential supports a superheavy freeze-in DM candidate with mass $\gtrsim 10^{10}\,\text{GeV}$ and requires a reheating temperature $\gtrsim 10^{12}\,\text{GeV}$, whereas the axion-like potential can yield a multi-TeV DM candidate for a dark freeze-out scenario, or a self-interacting GeV-scale DM candidate for an asymmetric DM model. The latter needs extension of the parameters, but can be accommodated naturally. They remain consistent with the combined constraints from DESI+CMB+DESY5 datasets within the $1\sigma$ level, along with the observed ICB angle. For the axion-type potential that emerges from a dark QCD sector, it is conceivable that there is an associated first-order phase transition in the DS. We have outlined that the dark freeze-out scenario may have associated gravitational waves with peak frequency in the range relevant for LISA Caprini:2019egz , TianQin TianQin:2020hid , and Taiji Hu:2017mde , while the asymmetric DM scenario suggests gravitational waves in the nano-Hz regime. A full analysis of the phase transition aspect of this model remains to be studied.
In conclusion, our framework offers a unified and minimal approach that simultaneously addresses the DESI anomaly, the origin of cosmic birefringence, and the production of particle DM consistent with the observed relic abundance measured by Planck. We highlight that upcoming DESI observations, especially with improved sensitivity at high redshifts, can provide critical tests of, or potentially falsify, this scenario.
@@ Acknowledgements.
SG acknowledges fruitful discussions with Yuichiro Nakai, Yu-Cheng Qiu, and Robert Shrock.

### v2

@@ 3.1 Dark matter in polynomial potential
$\Gamma=\frac{T}{8\,\pi^{4}}\,\int_{4m_{\psi}^{2}}^{\infty}ds\,s^{3/2}\,\sigma\left(s\right)_{\gamma\gamma\to\psi\psi}\,K_{1}\left(\frac{\sqrt{s}}{T}\right)\,$, (34)
@@
where we consider $n_{\rm eq}^{\rm DM}(T)=\frac{T}{2\pi^{2}}\,m_{\psi}^{2}\,K_{1}\left(m_{\psi}/T\right)$ to be the equilibrium DM number density, such that it provides a conservative limit on $T_{\text{rh}}$. For our choice of parameters, Eq. (41) is always satisfied. We emphasize that, for the benchmarks shown in Table 1, the DM–DE interaction modifies the DM mass at an earlier cosmological time at the level of $\sim 3\%$ for BP1, while its $\lesssim 1\%$ for the other benchmarks. To keep the final abundance unchanged, this variation can be reabsorbed in the unknown reheat temperature according to Eq. (39).
@@ 3.2 Dark matter in axion-type potential
    Dark freeze-out: The dark baryons $\psi$ can undergo freeze-out within the DS via, for example, $\bar{\psi}\psi\to\pi_{\rm D}\,\pi_{\rm D}$, where $\pi_{\rm D}$ are the dark pions. The dark quarks $Q,\bar{Q},{\mathbf{u,d}}$ can remain in equilibrium with each other via the mediation of dark gluons. Therefore, a notion of dark temperature $T_{\rm D}$ can be attributed. Despite $Q,\bar{Q}$ carrying hypercharge quantum numbers, and in the presence of dimension-5 EM coupling and effective operators relevant for $Q,\bar{Q}$ decay as mentioned in subsection 2.3, it is easy to ensure that the DS never comes in thermal equilibrium with the visible sector in the entire thermal cosmological history due to very large $f$. As long as one satisfies $T_{\rm rh}<f$, the heavy dark quarks $Q,\bar{Q}$ never become massless in the cosmological evolution. For example, consider the scattering $\gamma Q\to\gamma Q$ in the early Universe. It is sufficient to satisfy that the scattering rate for this process is less than the Hubble rate at the reheating temperature. This is satisfied when
@@
    obtained by comparing the interaction rate $n_{\rm eq}\,\langle\sigma\cdot v\rangle$ with the Hubble rate in a radiation dominated background, where we consider $\langle\sigma\cdot v\rangle\simeq C_{Y}^{4}\,\pi\,\alpha_{1}^{2}/m_{Q}^{2}$ and $n_{\rm eq}=\left(2\,\zeta(3)/\pi^{2}\right)\,T^{3}$ is the equilibrium number density of photons. The condition in Eq. (42) is easily satisfied with the chosen benchmark parameters in subsection 2.3. Similarly, we have checked that it is straightforward to satisfy the out-of-equilibrium condition for the dimension-5 EM coupling and the effective operators relevant for $Q,\bar{Q}$ decay. Furthermore, if one considers the operator $\langle\Phi\rangle^{*}/\Lambda_{d}^{3}\bar{\mathbf{u}}_{R}Qe_{R}e_{R}$, then the DS never comes in thermal equilibrium with the SM, provided
@@
    This points towards the dark QCD with dynamical scale $\Lambda_{\rm dQCD}\sim 10$ TeV, and therefore, the DM mass is $m_{\psi}\sim{\cal O}(50)$ TeV for $N_{D}=3$ as determined from Eq. (29). In this case, the DM behaves collisionless even at the dwarf galaxy scale as can be seen from Eq. (27).
@@
    Let us now turn to the question of stable DS relics other than the DM. For the current model, there are three dark pions $\pi_{\rm D}$, with mass $m_{\pi_{\rm D}}\simeq\sqrt{m_{\mathbf{q}}\Lambda_{\rm dQCD}}$, and can be determined from Eq. (26) for a given $\Lambda_{\rm dQCD}$ and $\varepsilon$. Clearly, for $\Lambda\simeq 2$ meV, as required by the DE energy density, and $\Lambda_{\rm dQCD}\simeq 10$ TeV, as required by the dark freeze-out mechanism, $m_{\pi_{\rm D}}\lesssim 10^{-14}$ eV for $\varepsilon\gtrsim 10^{-10}$. Further, in the absence of dark electromagnetism and weak interaction, the pions are stable. They also do not decay to the DE axion $\varphi$ due to very suppressed mixing $\simeq\Lambda_{\rm dQCD}/(4\pi f)\ll 1$. Hence, they behave as stable dark radiation and can be constrained by extra number of relativistic degrees of freedom $\Delta N_{\rm eff}$. If one assumes that the entropy is separately conserved in the visible and in the DS, then it is possible to write, $\xi\propto\left[g_{\star s}(T)/g_{\star s}^{\rm D}(T_{\rm D})\right]^{1/3}$, where $g_{\star s}$ and $g_{\star s}^{\rm D}$ are the number of relativistic degrees of freedom for entropy in the visible and the DS, respectively. As discussed in Cheung:2010gj , it is possible to constrain $\xi$ from the measurement of $\Delta N_{\text{eff}}=\frac{4}{7}\,\left(\frac{11}{4}\right)^{4/3}\,g_{\star}^{\rm D}(T_{\rm CMB})\,\xi(T_{\rm CMB})^{4}$, that takes care of the number of extra relativistic degrees of freedom beyond the SM at the time of CMB. Considering $\Delta N_{\text{eff}}\lesssim 0.34$ Planck:2018vyg from CMB observation, one finds, $\xi(T_{\rm CMB})\lesssim 0.48$, for $g_{\star}^{D}(T_{\rm CMB})=3$, as relevant for the three dark pions. Therefore, it is possible to tune $\Delta N_{\rm eff}$ by choosing an appropriate $\xi(T_{\rm CMB})$, consistent with the observations [^5] An enhanced number of relativistic species (e.g., dark radiation) increases the early-Universe radiation density and expansion rate, advancing matter-radiation equality and reducing the sound horizon. This leads to a larger inferred Hubble constant, with one extra species raising $H_{0}$ by about 7 km s-1 Mpc-1 in a flat Universe, enough to ease the Hubble tension Riess:2016jrr . The tension is effectively resolved for $\Delta N_{\text{eff}}\sim 0.4$, though it reappears once BAO data is included Bernal:2016gxb . A detailed analysis of possible non-zero $\Delta N_{\text{eff}}$ and its effect on the $H_{0}$, and $S_{8}$ tension, is beyond the scope of the present study.. It is also possible to introduce portal operators for the dark pions to decay to SM particles, and then both the visible and DS can be in thermal equilibrium, and can introduce laboratory probe channels for the DM Fujikura:2024jto . We do not pursue this possibility here. In passing, we mention that in the present scenario, it is possible to have dark glueballs generated in the thermal bath. However, they will be unstable and decay into the dark pions, given that the dark pion mass is tiny. Further, if the confining phase transition of this dark QCD sector is first-order, which is expected for large $N_{\rm D}$, it could be associated with gravitational waves (GWs). It is expected that the scale $\Lambda_{\rm dQCD}$ would be responsible for setting the reheating temperature scale after this phase transition. Therefore, the gravitational waves would have a peak frequency in the range Santos:2022hlx ; Fujikura:2023lkn
@@ 4 Conclusions
In this work, we have explored a consistent and minimal framework for an interacting DE-DM scenario that successfully explains the phantom crossing behavior of the evolving DE equation-of-state parameter, as indicated by recent DESI observations, without inducing any pathological features. We have introduced a dimension-5 effective operator involving the DE scalar and the SM photons, which accounts for the observed non-zero ICB angle. To explain the evolving DE behavior, we consider two well-motivated potentials for the DE scalar field: a polynomial potential, representative of tracker behavior in conventional quintessence models, and an axion-like potential naturally arising in strongly-coupled QCD-like dark sectors. We have provided a UV complete model for the latter. Furthermore, we have demonstrated how the DM can minimally obtain its relic density. Interestingly, the polynomial potential supports a superheavy freeze-in DM candidate with mass $\gtrsim 10^{10}\,\text{GeV}$ and requires a reheating temperature $\gtrsim 10^{12}\,\text{GeV}$, whereas the axion-like potential can yield a multi-TeV DM candidate for a dark freeze-out scenario, or a self-interacting GeV-scale DM candidate for an asymmetric DM model. The latter needs extension of the parameters, but can be accommodated naturally. They remain consistent with the combined constraints from DESI+CMB+DESY5 datasets within the $1\sigma$ level, along with the observed ICB angle. For the axion-type potential that emerges from a dark QCD sector, it is conceivable that there is an associated first-order phase transition in the DS. We have outlined that the dark freeze-out scenario may have associated gravitational waves with peak frequency in the range relevant for LISA Caprini:2019egz , TianQin TianQin:2020hid , and Taiji Hu:2017mde , while the asymmetric DM scenario suggests gravitational waves in the nano-Hz regime. A full analysis of the phase transition aspect of this model remains to be studied. In conclusion, our framework offers a unified and minimal approach that simultaneously addresses the DESI anomaly, the origin of cosmic birefringence, and the production of particle DM consistent with the observed relic abundance measured by Planck. We highlight that upcoming DESI observations, especially with improved sensitivity at high redshifts, can provide critical tests of, or potentially falsify, this scenario.
@@ Acknowledgements.
SG acknowledges fruitful discussions with Yuichiro Nakai, Yu-Cheng Qiu, Robert Shrock, and Hong-Yi Zhang.